# Masking Augmentation for Supervised Learning

### v2

@@ 4.5 Transfer learning
Improvement in pretraining can boost the performance of downstream tasks [29]. We measure the transfer learning performance of MaskSub using 800 epochs pretrained weight from Table 1. CIFAR-10 [31], CIFAR-100 [31], Oxford Flowers-102 [37], Stanford Cars [30] and iNaturalist [49] are used for finetuning datasets. We use the AdamW training recipe [47] and also evaluate performance when MaskSub (50%) is applied to the finetuning process. Table 10 shows the results. The backbone pretrained with MaskSub consistently outperforms the DeiT-III backbone across all cases. Moreover, when MaskSub is applied to the finetuning process, it further boosts performance in most cases except CIFAR [31].
4.6 Beyond image classification
Masking augmentation like MIM originates from a language training method [12]. It implies that MaskSub might be applicable to language domains similar to computer vision tasks. To verify general applicability, we apply MaskSub to two additional tasks beyond image classification: GLUE [50] benchmark and CLIP [41] pretraining. The first task is a text-classification benchmark GLUE [50]. We use BERT [12] as a pretrained model and apply MaskSub with 15% masking following the masking ratio of BERT. As shown in Table 12, MaskSub improves text-classification performance. MaskSub is also applied to CLIP [41] pretraining. Table 12 shows the results. CLIP trained with MaskSub (50%) shows improved zero-shot performance. Experimental details are in Section B. These results verify that MaskSub has significant impacts not only vision domain but also language and vision&language domain.
4.7 Extending to drop regularizations
@@
Computational costs ×$\times 1.0$$\times 1.5$$\times 2.0$$\times 2.0$
Table 13: Comparison of MaskSub variants. We validate drop-based variants of MaskSub: DropSub, and PathSub.
In Section 3.3, we propose MaskSub variants expanded to drop-related regularizations [44, 27]. We validate the performance of MaskSub variants on a 400 epochs training with Deit-III. We use masking [21] (50%), dropout [44] (0.2), and drop-path [27] (baseline + 0.1) for MaskSub, DropSub, and PathSub, respectively. Table 13 shows the results. Variants of MaskSub outperform the baseline. Among the three, MaskSub shows the best performance. Also, MaskSub has the lowest computation costs due to MAE [21]-style computation reduction. Thus, we conclude that MaskSub (50%) is the best in practice compared to variants with drop regularizations. Additional results are reported in Table A.5.
@@ 5 Conclusion
In this work, we have presented a new way to introduce masking augmentation to supervised learning. Our method, Masked Sub-model (MaskSub), is designed to leverage masking augmentation within a sub-model, which is separated from main training and uses a relaxed loss function. Our extensive analysis reveals that MaskSub effectively mitigates malicious effects of heavy masking while accelerating the convergence speed, yielding superior performance. We verify MaskSub on various training recipes, including diverse architecture. Notably, MaskSub demonstrates significant performance improvements across diverse scenarios. We claim that MaskSub is a substantial advancement in training recipes and contributes to building novel augmentation strategies.
@@ A.1 Training budget
We report an extensive comparison of the training budget in addition to Table 7. Table A.1 shows the performance improvements of MaskSub compared to training recipes with increased epochs to have the same training costs as MaskSub. In most cases, MaskSub outperforms its counterparts even with the same training computation costs. The results imply that MaskSub is a better solution than increasing training epochs to improve model performance. Note that the performance of MAE finetuning is lower than the original epoch since it utilizes benefits from early stop training to prevent overfitting.
@@
Table A.1: Comparison in the same training budget. All training has been conducted with NVIDIA V100 8 GPUs. GPU days refer to the number of days required for training when using a single V100 GPU.
@@ A.2 Downstream tasks
Semantic segmentation. Following the training setup of the ADE-20k sementaic segmentation (provided in BEiT v2 [38]), we train UpperNet [57] with the ViT backbones on ADE-20k [67]. Table A.3 shows the results of single-scale and multi-scale evaluations. On both evaluations, the backbone pretrained with MaskSub demonstrates superior performance, consistent for the ViT-B and ViT-L backbones.
@@
Table A.2: Semantic segmentation on ADE-20k. UpperNet for ViT backbone is trained with the BEiTv2 segmentation recipe.
Single-scale mIoU Multi-scale mIoU
DeiT-III + MaskSub DeiT-III + MaskSub
ViT-B 48.8 49.4 (+0.6) 49.7 50.2 (+0.5)
ViT-L 51.7 52.2 (+0.5) 52.3 52.7 (+0.4)
@@
Table A.2: Semantic segmentation on ADE-20k. UpperNet for ViT backbone is trained with the BEiTv2 segmentation recipe.
@@ A.3 Robustness
Model +MaskSub IN-1k IN-V2 IN-Real IN-A IN-R ObjNet SI-size SI-loc SI-rot
ViT-S - 81.4 70.1 87.0 23.4 46.4 32.6 55.0 39.8 37.8
81.7 71.0 87.4 26.9 47.2 33.5 56.7 42.5 39.9
ViT-B - 83.8 73.4 88.2 36.8 54.1 35.7 58.0 42.7 41.5
84.2 74.0 88.6 41.9 54.4 37.2 59.0 44.8 43.3
ViT-L - 84.9 74.8 88.8 45.3 57.4 38.8 59.8 46.5 45.0
85.3 75.8 89.2 51.1 58.5 40.0 60.2 46.8 45.9
ViT-H - 85.2 75.7 89.2 51.9 58.8 40.1 61.9 49.0 46.8
85.7 76.5 89.6 58.3 59.9 41.7 62.4 50.1 48.4
Table A.4: Robustness benchmark. Table shows the robustness benchmark for ViT pretrained with/without MaskSub. In all metrics, higher scores indicate better results.
@@ A.4 Extending to drop regularizations
To generalize MaskSub to drop regularizations, we demonstrate MaskSub with drop-out [44] (DropSub) and drop-path [27] (PathSub) in Table 13. In this section, we present additional analysis to verify MaskSub variants. The experiment setup is the same as in Section 3.2. The results are shown in Table A.5. We validate three MaskSub methods for three kinds of regularizations: random masking [21], drop-out [44], and drop-path [27]. “Original” means the original training recipe trained with Eq. 1, and none of the drop regularizations is used. “Single model” shows the performance when the network is trained with Eq. 2. Note that it is a common practice to use drop-based techniques. We compare those common practices with “MaskSub variants”, models trained with Eq. 1 and 3. For analysis, we measured Eq. 1 “train loss (original)” and Eq. 2 “train loss (drop)” regardless of the loss used for training. It shows how losses changed by training setting. The results demonstrate that MaskSub improves training in all three cases. In all cases, MaskSub improves original and drop loss convergence, which is connected to superior accuracy compared to original training.
@@ A.5 Ablation studies
We conduct ablation studies of MaskSub using DeiT-III training settings to validate the role of KD loss and the effects of hyperparameters. Table A.6 shows the impacts of KD loss in MaskSub. Performance improvement by MaskSub is significantly limited when KD loss is not used. Analysis for masking ratio and loss weights are shown in Table A.8 and Table A.8 respectively. Note that we use DeiT-III 100 epochs setting and baseline without MaskSub is 77.4.
@@
Mask ratio 0% 25% 40% 50% 60% 75% Accuracy 78.4 79.0 79.2 79.2 79.3 79.2 Costs ×$\times 2.0$$\times 1.75$$\times 1.6$$\times 1.5$$\times 1.4$$\times 1.25$
Weights (CE, KD) (1, 0) ($\frac{3}{4}$, $\frac{1}{4}$) ($\frac{2}{3}$, $\frac{1}{3}$) ($\frac{1}{2}$, $\frac{1}{2}$) ($\frac{1}{3}$, $\frac{2}{3}$) ($\frac{1}{4}$, $\frac{3}{4}$) Accuracy 77.4 78.5 78.7 79.2 79.2 78.5
@@ A.6 Reporting mean and standard deviation
Mean and standard deviation values for transfer learning in Table 10 are shown in Table A.9. MaskSub demonstrates significant gains, which surpass the standard deviation of performance. Table A.10 presents short training (300 epochs) results. MaskSub shows substantial improvements when it is applied to both pretraining and finetuning processes.
Table A.12 shows 400 epochs training with DeiT-III [47], which is reported in Table 13 of the paper. MaskSub variants improve the performance of ViT training. Compared to the variants, MaskSub demonstrates the best performance.
Table A.11 shows the mean and standard deviation values for Table A.5. Table A.11 shows the superiority of our MaskSub in additional regularization training.
@@
Table A.9: Mean and std for transfer learning to small scale datasets. Table shows ‘mean $\pm$ std’ values for transfer learning performance with/without MaskSub. We measure the performance when MaskSub is applied to pretraining and finetuning.
@@
Table A.10: Transfer learning at short (300 epochs) training. Table shows ‘mean $\pm$ std’ values for transfer learning at 300 epochs. We measure the performance when MaskSub is applied to pretraining and finetuning.
@@ Appendix B Implementation details
Most experiments in the paper were performed on a machine with NVIDIA V100 8 GPUs. The exceptions were DeiT-III [47] experiments for ViT-L and ViT-H in Table 1 and object detection in Table A.3, conducted with NVIDIA A100 64 GPUs. Also, we use a single NVIDIA V100 for transfer learning in Table 10. We strictly follow original training recipes for experiments. We denote details of the training recipes to clarify our implementation details and assist in reproducing our results.
Image Classification. Table B.1 shows training recipes used for Table 1, 3, 5, 8, 9, and 13 of the paper. It demonstrates that our MaskSub is validated on various training recipes that cover diverse regularization and optimizer settings and achieves consistent improvement on all settings, which exhibits the general applicability of MaskSub. Note that MaskSub is applied to all recipes with the same masking ratio of 50%. Thus, MaskSub does not require hyper-parameter tuning specialized for each recipe. Model-specific training recipes of DeiT-III [47] are reported in Table B.2. DeiT-III achieves strong performance with sophistically tuned training parameters mainly focused on input size and drop-path rate. It makes improving DeiT-III more challenging than other recipes, yet our MaskSub accomplishes this with a significant performance gap.
@@
Object detection and instance segmentation We use Detectron2 [56] on the COCO [35] dataset. Among various recipes in the Detectron2 library, we use ViTDet [33], which is a recent and strong recipe for a ViT-based detector. We train ViTDet Cascaded Mask-RCNN with DeiT-III and DeiT-III + MaskSub pretrained backbone and report performance after MSCOCO 100 epoch training.
@@
Table B.2: Model specific recipes of DeiT-III [47]. The table shows model size and training length-specific training arguments used for the DeiT-III recipe. In addition to Table B.1, DeiT-III utilizes drop-path and image size to adjust the recipe for diverse model sizes and training lengths.

### v3

@@ 4.6 Transfer learning
Improvement in pretraining can boost the performance of downstream tasks [31]. We measure the transfer learning performance of MaskSub using 800 epochs pretrained weight from Table 1. CIFAR-10 [33], CIFAR-100 [33], Oxford Flowers-102 [40], Stanford Cars [32] and iNaturalist [52] are used for finetuning datasets. We use the AdamW training recipe [50] and also evaluate performance when MaskSub (50%) is applied to the finetuning process. Table 7 shows the results. The backbone pretrained with MaskSub consistently outperforms the DeiT-III backbone across all cases. Moreover, when MaskSub is applied to the finetuning, it further boosts performance except CIFAR [33].
We verify transfer learning to semantic segmentation task on ADE-20k [71]. We train UperNet [60] training recipe [41] and utilize pretrained weight from Table 1. Table 10 shows the segmentation results of single-scale and multi-scale evaluations. On both evaluations, the backbone pretrained with MaskSub demonstrates superior performance, consistent for ViT-B and ViT-L.
4.7 Beyond vision domain
MaskSub can be extended to domains beyond images, as long as the masking is applicable. Thus, we apply MaskSub to two additional tasks beyond the image domain: GLUE [53] benchmark and CLIP [44] pretraining. The first task is a text-classification benchmark GLUE [53]. We use BERT [13] as a pretrained model and apply MaskSub with 15% masking following the masking ratio of BERT. As shown in Table 11, MaskSub improves text-classification performance. MaskSub is also applied to CLIP [44] pretraining. Table 12 shows the results. CLIP trained with MaskSub (50%) shows improved zero-shot performance. Experimental details are in Appendix. These results verify that MaskSub has remarkable impacts not only on the vision but also on the language and vision&language domain.
@@ 4.7 Beyond vision domain
Computation ×$\times 1.0$$\times 1.5$$\times 2.0$$\times 2.0$
Table 13: Comparison of MaskSub variants. We validate drop-based variants of MaskSub. The sub-branch training improves performance with other drop-methods. But, MaskSub shows the best improvement with the smallest computations.
4.8 Extending to drop regularizations
In Section 3.3, we expand MaskSub with drop regularizations [47, 28]. We validate the performance of MaskSub variants on a 400 epochs training with Deit-III. We use masking [22] (50%), dropout [47] (0.2), and drop-path [28] (baseline + 0.1) for MaskSub, DropSub, and PathSub, respectively. Table 13 shows the results. Variants of MaskSub outperform the baseline. Among the three, MaskSub shows the best performance. Also, MaskSub has the lowest computation costs due to MAE [22]-style computation reduction. Thus, we conclude that MaskSub (50%) is the best in practice compared to variants with drop regularizations. Note that Table A.5 in Appendix includes more results.
@@ 5 Conclusion
In this work, we have presented a new way to introduce masking augmentation to supervised learning. Our method, Masked Sub-branch (MaskSub), is designed to leverage masking augmentation within a sub-branch, which is separated from main training and uses a relaxed loss function. Our extensive analysis reveals that MaskSub effectively mitigates malicious effects of heavy masking while accelerating the convergence, yielding superior performance. We verify MaskSub on various training recipes with diverse architecture. Notably, MaskSub demonstrates impressive performance improvements across various scenarios. We claim that MaskSub is a substantial advancement in training recipes and contributes to using augmentations.
@@ A.1 Training budget
We report an extensive comparison of the training budget in addition to Table 2 and 8. Table A.1 shows the performance improvements of MaskSub compared to training recipes with increased epochs to have the same training costs as MaskSub. In most cases, MaskSub outperforms its counterparts even with the same training computation costs. The results imply that MaskSub is a better solution than increasing training epochs to improve model performance. Note that the performance of MAE finetuning is lower than that of the original epoch since it utilizes the benefits of early stop training to prevent overfitting. Table A.2 presents a comparison with other training methods using additional augmentations in MAE [22] fine-tuning task. Similar to Table 8, MaskSub shows the impressive trade-off between performance and computation costs, even in the fine-tuning task. Fig. A.1 shows training analysis in Fig. 2 with the same training budget setting. We use GPU days as an x-axis of analysis. MaskSub is effective for loss convergence and accuracy, even considering the additional training budget.
Table A.1: Comparison under the equivalent training budget. All trainings are conducted on NVIDIA V100 8 GPUs. GPU days refer to the number of days required for training when using a single V100 GPU.
@@
(a) Accuracy
(b) Train loss (standard)
(c) Train loss (masking)
(d) Gradient norm
Figure A.1: MaskSub training analysis for training budget. The figures show training analysis experiments in Fig. 2 with the training budget (GPU days) as x-axis. Even considering its additional training budgets, MaskSub effectively improves convergence and accuracy.
Table A.2: MAE finetuning comparison. The table displays performances and GPU costs for MAE ViT-B fine-tuning.
Method Accuracy GPU days
Baseline [50] 83.6 6.0
DataAug [26] 83.1 (-0.5) 11.8 (+97%)
GradAug [64] 84.0(+0.4) 23.4 (+290%)
CoSub [51] 83.8 (+0.2) 11.0 (+83%)
MaskSub 83.9 (+0.3) 8.6 (+43%)
@@ A.2 Downstream tasks
Transfer learning performances are presented in Table 7 in the main paper.
Semantic segmentation results are included in Table 10 in the main paper.
@@ A.3 Robustness
Table A.4: Robustness benchmark. Table shows the robustness benchmark for ViT pretrained with/without MaskSub. $\uparrow$ means higher score is better robustness, while $\downarrow$ indicates lower score is better.
Model +MaskSub IN-1k($\uparrow$) IN-V2($\uparrow$) IN-Real($\uparrow$) IN-A($\uparrow$) IN-R($\uparrow$) IN-C($\downarrow$) ObjNet($\uparrow$) SI-size($\uparrow$) SI-loc($\uparrow$) SI-rot($\uparrow$)
ViT-S - 81.4 70.1 87.0 23.4 46.4 44.8 32.6 55.0 39.8 37.8
81.7 71.0 87.4 26.9 47.2 43.6 33.5 56.7 42.5 39.9
ViT-B - 83.8 73.4 88.2 36.8 54.1 38.1 35.7 58.0 42.7 41.5
84.2 74.0 88.6 41.9 54.4 36.9 37.2 59.0 44.8 43.3
ViT-L - 84.9 74.8 88.8 45.3 57.4 33.9 38.8 59.8 46.5 45.0
85.3 75.8 89.2 51.1 58.5 32.4 40.0 60.2 46.8 45.9
ViT-H - 85.2 75.7 89.2 51.9 58.8 32.8 40.1 61.9 49.0 46.8
85.7 76.5 89.6 58.3 59.9 31.2 41.7 62.4 50.1 48.4
@@ A.4 Extending to drop regularizations
To generalize MaskSub to drop regularizations, we demonstrate MaskSub with drop-out [47] (DropSub) and drop-path [28] (PathSub) in Table 13. In this section, we present additional analysis to verify MaskSub variants. The experiment setup is the same as in Section 3.2. The results are shown in Table A.5. We validate three MaskSub methods for three kinds of regularizations: random masking [22], drop-out [47], and drop-path [28]. “Original” means the original training recipe trained with Eq. 1, and none of the drop regularizations is used. “Single model” shows the performance when the network is trained with Eq. 2. We compare those common practices with “MaskSub variants”, models trained with Eq. 1 and 3. For analysis, we measured Eq. 1 “train loss (original)” and Eq. 2 “train loss (drop)” regardless of the loss used for training. It shows how losses changed by training setting. The results demonstrate that MaskSub improves training in all three cases. In all cases, MaskSub improves original and drop loss convergence, which is connected to superior accuracy compared to original training.
@@ A.5 Ablation studies
We conduct ablation studies of MaskSub using DeiT-III training settings to validate the role of KD loss and the effects of hyperparameters. Table A.6 shows the impacts of KD loss in MaskSub. Performance improvement by MaskSub is significantly limited when KD loss is not used. Table A.7 and Table A.8 show analysis for masking ratio and loss weights, respectively. Note that we train for 100 epochs with the DeiT-III setting, and the baseline without MaskSub is 77.4. In Table A.9, we test additional ways to improve MaskSub’s performance: multiple sub-branch and photometric augmentation on sub-branch. We train ViT-B/16 with 400 epochs DeiT-III setting with MaskSub variants. Single sub-branch shows the best performance with the smallest computation budget. Color-jittering improves MaskSub with negligible cost increases, but Gaussian blur is ineffective.
@@
Mask ratio 0% 25% 40% 50% 60% 75%
Accuracy 78.4 79.0 79.2 79.2 79.3 79.2
Costs ×$\times 2.0$$\times 1.75$$\times 1.6$$\times 1.5$$\times 1.4$$\times 1.25$
@@
Weights (CE, KD)
(1, 0) ($\frac{3}{4}$, $\frac{1}{4}$) ($\frac{2}{3}$, $\frac{1}{3}$) ($\frac{1}{2}$, $\frac{1}{2}$) ($\frac{1}{3}$, $\frac{2}{3}$) ($\frac{1}{4}$, $\frac{3}{4}$)
Accuracy 77.4 78.5 78.7 79.2 79.2 78.5
Table A.9: Multiple sub-branch and photometric aug. We test multiple sub-branch settings and additional photometric augmentation on sub-branch to further improve the performance. The result shows that a single sub-branch is the best for accuracy and computation. Additional color jittering on the sub-branch improves performance, but Gaussian blur is not effective.
MaskSub +
2-branch
masking
3-branch
masking
Color
jitter
Gaussian
blur
Accuracy 84.10 84.03(-0.07) 83.76(-0.34) 84.18(+0.08) 84.05(-0.05)
Costs $\times$1.5 $\times$2.0 $\times$2.5 $\times$1.5 $\times$1.5
@@ A.6 Reporting mean and standard deviation
Mean and standard deviation values for transfer learning in Table 7 are shown in Table A.10. MaskSub demonstrates meaningful performance gains, which surpass the standard deviation of performance. Table A.11 presents short training (300 epochs) results. MaskSub shows substantial improvements when applied to pretraining and finetuning processes. Table A.13 shows 400 epochs training with DeiT-III [50], which is reported in Table 13 of the paper. MaskSub variants improve the performance of ViT training. Compared to the variants, MaskSub demonstrates the best performance. Table A.12 shows the mean and standard deviation values for Table A.5. Table A.12 shows the superiority of our MaskSub in additional regularization training.
Table A.10: Mean and std for transfer learning to small scale datasets. The table shows ‘mean $\pm$ std’ values for transfer learning performance with/without MaskSub. We measure the performance when MaskSub is applied to pretraining and finetuning.
@@
Table A.11: Transfer learning at short (300 epochs) training. The table shows ‘mean $\pm$ std’ values for transfer learning at 300 epochs. We measure the performance when MaskSub is applied to pretraining and finetuning.
@@ Appendix B Implementation details
Most experiments in the paper were performed on a machine with NVIDIA V100 8 GPUs. The exceptions were DeiT-III [50] experiments for ViT-L and ViT-H in Table 1 and object detection in Table A.3, conducted with NVIDIA A100 64 GPUs. Also, we use a single NVIDIA V100 for transfer learning, as shown in Table 7. We strictly follow original training recipes for experiments. We denote details of the training recipes to clarify our implementation details and assist in reproducing our results.
Image Classification. Table B.1 shows training recipes used for Table 1, 3, 5, 9, 13 and A.2. It demonstrates that our MaskSub is validated on various training recipes that cover diverse regularization and optimizer settings and achieves consistent improvement on all settings, which exhibits the general applicability of MaskSub. Note that MaskSub is applied to all recipes with the same masking ratio of 50%. Thus, MaskSub does not require hyper-parameter tuning specialized for each recipe. Model-specific training recipes of DeiT-III [50] are reported in Table B.2. DeiT-III achieves strong performance with sophistically tuned training parameters mainly focused on input size and drop-path rate. It makes improving DeiT-III more challenging than other recipes, yet our MaskSub accomplishes this with a reasonable performance gap.
@@
Object detection and instance segmentation We use Detectron2 [59] on the COCO [37] dataset. Among various recipes in the Detectron2 library, we use ViTDet [35], which is a recent and strong recipe for a ViT-based detector. We train ViTDet Cascaded Mask-RCNN with DeiT-III and DeiT-III + MaskSub pretrained backbone and report performance after MSCOCO 100 epochs training.
@@
Table B.2: Model specific recipes of DeiT-III [50]. The table shows the model size and training length-specific training arguments used for the DeiT-III recipe. In addition to Table B.1, DeiT-III utilizes drop-path and image size to adjust the recipe.